OpenStudy (vshiroky):

Find an equation of a line that has the given slope and passes through the given point. m=-6 (4,5)

OpenStudy (vshiroky):

Not sure if I did this right... y-y1=-6(x-x1) y=-6x+29

ganeshie8 (ganeshie8):

you can check if the line is correct or not, by plugging in the given point (4, 5)

ganeshie8 (ganeshie8):

y=-6x+29 plugin (4, 5)

OpenStudy (vshiroky):

Looks like it is to me.

ganeshie8 (ganeshie8):

y=-6x+29 5 = -6*4 + 29 5 = -24 + 29 5 = 5

ganeshie8 (ganeshie8):

So yes !! good job :)